OpenStudy (goldie1962):

Factor the algebraic expression below in terms of a single trigonometric function. cos x - sin ^2x - 1

OpenStudy (danjs):

you remember the property from the unit circle. \[\large \sin^2(x)+\cos^2(x) = 1\]

OpenStudy (goldie1962):

Yes I do

OpenStudy (danjs):

Solve that for sin^2(x) \[\sin^2(x)=1-\cos^2(x)\] Put that sin^2(x) into the expression for sin^2(x) \[\cos(x)-(1-\cos^2(x)) -1\]

OpenStudy (danjs):

That gives you \[\large \cos^2(x) + \cos(x) - 2\]

OpenStudy (danjs):

You can factor that out to \[\large (\cos x + 2)*(\cos x - 1)\]

OpenStudy (goldie1962):

Oh wow okay thanks!